Write down equations for the following constraints of mechanical systems
and classify them with respect to whether they are scleronomic or rheonomic and
holonomic or not nonholonomic.
a) A small ball which rolls down on the surface of a big ball without friction.
b) The suspension point of a planar pendulum carries out harmonic horizontal
oscillations.
c) A bead moving freely on a long straight wire through the origin, which rotates
in the x-y plane with constant angular velocity ω about the z-axis.
d) A bead slides along a straight wire. The wire has constant inclination α
towards the horizontal x-y-plane and moves with constant acceleration in x
direction.
1
Expert's answer
2020-11-23T10:28:31-0500

a) A small ball which rolls down on the surface of a big ball without friction- It is non-holonomic and it's equation will be "r^2-a^2>0"

b)The suspension point of a planar pendulum carries out harmonic horizontal oscillations, it is an example of scleronomic system."|\overrightarrow{r}^2|=l^2"

c)A bead moving freely on a long straight wire through the origin, which rotates

in the x-y plane with constant angular velocity ω about the z-axis. It is an example of a non-holonomic constraint."z_1=z_2=0"

"(x_1-x_2)^2+(y_1-y_2)^2=l^2"

d)A bead slides along a straight wire. The wire has constant inclination α

towards the horizontal x-y-plane and moves with constant acceleration in x

direction. It is a example of non-holnomic. and it's equation will be "r^2-a^2>0"